# Self-Augmented Robot Trajectory: Efficient Imitation Learning via Safe Self-augmentation with Demonstrator-annotated Precision

### v1

@@ 1 Introduction
Despite these advantages, many existing IL methods depend on extensive datasets of human demonstrations, the collection of which is both labor intensive and time consuming [2, 3]. To mitigate this bottleneck, recent efforts have explored robotic data augmentation, that is, automatically expanding the training set by allowing robots to explore and generate supplementary data [4, 5, 6, 7]. Although this strategy can effectively reduce the need for human involvement, methods based on random exploration are ill-suited to clearance-limited tasks, where uninformed movements often lead to collisions and necessitate manual environment resets, as illustrated in Fig. 1. As a result, the overall human effort may not be significantly reduced. Addressing these issues often requires additional assumptions regarding the environment or task structure, which limit the general applicability of IL.
To overcome these limitations, the self-augmented robot trajectory (SART) is proposed, which is a safe and efficient imitation learning framework that enables policy learning from a single human demonstration augmented by automatically generated data within explicitly annotated precision boundaries, as illustrated in Fig. 1. In SART, a human expert provides a single demonstration trajectory for the task and annotates safety regions—represented as spheres—at each key waypoint. These regions define the permissible operating space for the robot’s end effector, allowing autonomous augmentation, which is both safe and effective. Extensive evaluations in both simulation and real-world manipulation tasks demonstrate that SART achieves substantially higher success rates than baselines, while also reducing the overall human duration required for dataset collection.
@@ 5.1.1 Environments and Tasks
Refer to caption Peg-in-hole
Refer to caption Door opening
@@ 5.2.2 Which design of SART makes the best performance?
Overall, the ablation results demonstrated position augmentation, merging with the original demonstration, and a progression-based augmentation design (rather than returning to sphere centers) are essential for SART’s performance gains. Although orientation augmentation is less critical in the current benchmarks, it provides robustness and supports the overall generality of the framework.
@@ 6 Discussion
A key limitation is that precision boundaries allow collision-free augmentation but do not explicitly address contact. Consequently, the application of SART in contact-rich tasks remains challenging. Future studies can integrate compliance-aware control and contact-informed augmentation strategies (e.g., MILES-style formulations) to enable safe and effective exploration in contact-rich scenarios [7].

### v2

@@ 1 Introduction
Despite these advantages, many existing IL methods depend on extensive datasets of human demonstrations, the collection of which is both labor intensive and time consuming [5, 9]. To mitigate this bottleneck, recent efforts have explored robotic data augmentation, that is, automatically expanding the training set by allowing robots to explore and generate supplementary data [7, 18, 1, 17]. Although this strategy can effectively reduce the need for human involvement, methods based on random exploration are ill-suited to clearance-limited tasks, where uninformed movements often lead to collisions and necessitate manual environment resets, as illustrated in Fig. 1. As a result, the overall human effort may not be significantly reduced. Addressing these issues often requires additional assumptions regarding the environment or task structure, which limit the general applicability of IL. Addressing these issues often requires additional environmental or structural assumptions, limiting the generality of IL. For example, Papagiannis et al. [17] reduce collisions by constraining exploration to a narrow, fixed region around each waypoint—effective for ensuring safety near contact-critical areas but overly restrictive in open spaces, leading to inefficient exploration. This motivates the development of a method that achieves a better balance between safety and efficiency across varying task conditions.
To this end, the self-augmented robot trajectory (SART) is proposed, which is a safe and efficient imitation learning framework that enables policy learning from a single human demonstration augmented by automatically generated data within explicitly annotated precision boundaries, as illustrated in Fig. 1. In SART, a human expert provides a single demonstration trajectory for the task and annotates safety regions—represented as spheres—at each key waypoint. These regions define the permissible operating space for the robot’s end effector, allowing autonomous augmentation, which is both safe and effective. Extensive evaluations in both simulation and real-world manipulation tasks demonstrate that SART achieves substantially higher success rates than baselines, while also reducing the overall human duration required for dataset collection.
@@ 5.1.1 Environments and Tasks
Peg-in-hole
Door opening
@@ 5.2.2 Which design of SART makes the best performance?
These ablation results demonstrated position augmentation, merging with the original demonstration, and a progression-based augmentation design (rather than returning to sphere centers) are essential for SART’s performance gains. Although orientation augmentation is less critical in the current benchmarks, it provides robustness and supports the overall generality of the framework. Further experiments were conducted to further analyze alternative convergence strategies and sampling schemes (Appendix C), as well as to compare SART with contact-free MILES variants using different fixed sphere sizes (Appendix D).
@@ 6 Discussion
Although SART demonstrated significantly better performance than behavioral cloning (BC) even when trained from a single demonstration, this result may partly stem from the use of a single trajectory that avoids the inherent ambiguity of human demonstrations. In IL, human demonstrations often represent multi-valued mappings, where multiple distinct trajectories can achieve the same goal. In this work, SART focuses on validating the efficacy of single-demonstration augmentation using a simple multilayer perceptron (MLP)-based policy; however, such a model is limited in its ability to represent multiple optimal strategies demonstrated by humans. Extending SART to augment and learn from a small set of diverse human demonstrations could further amplify its performance while addressing this multimodal challenge. To achieve this, future work should explore more expressive policy architectures, such as diffusion-based policies [4], which are capable of capturing multiple feasible action distributions and thus better leverage diverse human data.
@@
Beyond these task-specific extensions, a broader opportunity lies in scaling SART toward large-scale robot learning. Recent advances in robotic foundational models [5, 9] have shown that large, cross-task datasets leveraging broad pre-trained visuomotor representations can substantially reduce the number of demonstrations required for new tasks. While these models rely on massive and diverse datasets to achieve generalization, the proposed SART framework can be viewed as a complementary approach that supports scalable and safe data collection at the per-task level. By autonomously transforming a small number of human demonstrations into diverse, collision-free trajectories, SART has the potential to contribute to the development of scalable datasets beneficial for such models. As a potential extension, SART could be expanded to increase data diversity in simulation—e.g., by varying object sizes [11] or poses [26]—which may help support broader generalization in robot learning.
While SART enables safe data augmentation within annotated precision boundaries, the manual effort required to specify these regions may still impose an additional burden on the human operator. To further reduce this overhead, a promising extension would be to incorporate classical motion-planning approaches that automatically generate collision-free trajectories in contact-free regions [25, 21]. Such integration could complement SART’s safety guarantees by automating the exploration process without relying on detailed human annotations.
Despite the safety achieved in such contact-free augmentation, SART does not explicitly handle contact, and its applicability to contact-rich tasks remains limited. A future direction is to develop a unified framework that integrates the proposed contact-free SART with contact-aware augmentation methods [17], enabling safer, more reactive data generation in contact-rich scenarios.
@@ Appendix B Annotation Interface Implementation
Figure 9: Comparison of SART with ablated variants. Only simulation results are reported. Each result is averaged over 55 independently trained policies (with different random seeds) and $N=40$ training trajectories, where each policy is tested with $10$ rollouts, yielding $50$ executions per method. An asterisk () indicates cases where removing a SART design component leads to a statistically significant performance drop ($t$-test, $p<0.05$).
Appendix C Additional Results: SART Ablations
To further analyze the design factors that contribute to SART’s performance, additional ablation experiments were conducted to evaluate 1. the effect of alternative convergence strategies and 2. the impact of sampling positions inside the precision sphere rather than on its surface. These analyses clarify how SART’s convergence and sampling formulations affect safety and efficiency during data augmentation.
As shown in Fig. 9, several variants were tested:
  • Convergence to Next-Sphere Center (w/ Conv. to Next Sphere Cent.) The robot converges to the center of the next precision sphere $\boldsymbol{c}^{\mathrm{pos}}_{i+1}$ instead of the current convergence point $\boldsymbol{c}^{\mathrm{pos}}_{t^{*}}$. Because the annotated spheres are always positioned consecutively under the current annotator, this variant yields performance identical to the default SART. Note that this arrangement was not intentionally constrained.
  • Convergence to Next-Sphere’s Convergence Point (w/ Conv. of Next Sphere) The convergence waypoint is changed to the next sphere’s convergence point $\boldsymbol{\kappa}^{*}_{i+1}$ instead of the current sphere’s $\boldsymbol{\kappa}^{*}_{i}$. This modification largely degrades the smoothness of the trajectory, leading to unstable transitions and reduced policy performance.
  • Extended Convergence Point (w/ Extended Conv. Point) In this variant, the convergence point is set twice as far from the sphere center, i.e., $\left\|\mathbf{c}^{\mathrm{pos}}_{t}-\mathbf{c}^{\mathrm{pos}}_{i}\right\|_{2}^{2}>(2r_{i})^{2}$ in Eq. (6). This extension frequently induces collisions during self-augmentation, which leads to the inclusion of multiple failed trajectories in the training dataset and consequently causes a significant performance degradation.
  • Sampling from Inside the Sphere (w/ Sample Inside Sphere) Random poses $\boldsymbol{p}^{k}_{i}$ are sampled inside the precision sphere rather than on its surface, $\left\|\boldsymbol{p}^{k}_{i}-\boldsymbol{c}^{\mathrm{pos}}_{i}\right\|_{2}<r_{i}$. Because this reduces the spatial range of exploration, the resulting dataset lacks diversity, producing poorer generalization compared with the default configuration.
Overall, these results confirm that SART’s design—particularly the use of adaptive precision boundaries and convergence to the nearest feasible waypoint—plays a crucial role in balancing safety and efficiency during augmentation.
Figure 10: Comparison of SART with contact-free MILES variants using different fixed sphere sizes. Simulation results are reported. Each result is averaged over 55 independently trained policies (with different random seeds) and $N=40$ training trajectories, where each policy is tested with $10$ rollouts, yielding $50$ executions per method. An asterisk () indicates cases where a configuration yields statistically lower performance than SART ($t$-test, $p<0.05$).
Appendix D Additional Results: Comparison between SART and Contact-Free MILES Variants
To further analyze the effect of the augmentation boundary size, additional experiments were conducted comparing SART with contact-free MILES variants that used different fixed sphere radii: $1\,\mathrm{cm}$, $2\,\mathrm{cm}$ (default), and $10\,\mathrm{cm}$. These settings evaluate whether reducing or enlarging the augmentation region influences policy robustness in clearance-limited tasks.
As shown in Fig. 10, all contact-free MILES variants—regardless of sphere size—exhibited significantly poorer performance than SART. Increasing the fixed sphere radius in contact-free MILES from $1\,\mathrm{cm}$ to $10\,\mathrm{cm}$ led to a slight improvement in performance due to broader exploration, but the results remained substantially below those achieved by SART. When the sphere size was too small ($1\,\mathrm{cm}$), the robot explored an overly restricted region, resulting in limited data diversity and weaker generalization. Conversely, when the sphere size was excessively large ($10\,\mathrm{cm}$), most of the sampled poses fell outside the feasible collision-free region, leaving only a small portion of the augmentation space effectively usable. As a result, the generated dataset became sparse and unbalanced, leading to degraded task success. In contrast, SART adaptively defines precision boundaries according to task-specific spatial constraints, maintaining both collision safety and trajectory diversity. These results confirm that SART’s adaptive, boundary-aware formulation is substantially more effective and robust than globally fixed augmentation regions.